OpenStudy (anonymous):

What is the difference of r+2/r+4 - 3/r+1

mathslover (mathslover):

Is it : \(\cfrac{r+2}{r+4} - \cfrac{3}{r+1} \) ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Do you know how to take LCM ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@passive13

mathslover (mathslover):

Moomoocow123, can you tell me what you tried yet?

OpenStudy (anonymous):

r^2-10/(r+4)(r+1) is one of the options?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

think so

OpenStudy (anonymous):

@passive13, buddy we've gone over this...you cant give out an answer like that.

OpenStudy (jhannybean):

Multiply, \[\frac{ \color{red}{(r+1)} }{ \color{red}{(r+1)} }*\frac{ (r+2) }{ (r+4) }-\frac{ \color{blue}{(r+4)} }{ \color{blue}{(r+4)} }*\frac{ 3 }{ (r+1) }\]

OpenStudy (anonymous):

i shall start charging people then lol

OpenStudy (jhannybean):

\[\frac{ (r+1)(r+2) }{ (r+1)(r+4) }-\frac{ 3(r+4) }{ (r+4)(r+1) }\]\[\frac{ r^2+2r+r+2 }{ (r+1)(r+4) }- \frac{ 3r+12 }{ (r+1)(r+4) }\] combine.\[\frac{ r^2+3r+2-3r-12 }{ (r+1)(r+4) }\]

OpenStudy (jhannybean):

Simplify and you've got your answer.